# Stop treating ‘AGI’ as the north-star goal of AI research

### v3

@@ 2.5 Generality Debt
Relying on the generality or flexibility of tools to postpone crucial engineering, scientific, or societal decisions.
@@
This trap includes the appeal to many different notions of generality at play in machine learning: (1) variety of tasks (Hernandez-Orallo & O’hEigeartiagh, 2018); (2) capability to be trained for “any task” vs. ability to perform many predefined tasks (Hernandez-Orallo & O’hEigeartiagh, 2018); (3) whether the task or data distribution the model is being evaluated on is “seen” or “unseen” (i.e., available, or not, to the model during its training phase) (Altmeyer et al., 2024); (4) variety of data in model input/output, such as structured vs unstructured, modality, etc.; (5) whether the performance of the model reflects “performance considered ‘surprising’ to humans” (Altmeyer et al., 2024); (6) variety of goals; (7) ability to “accept a general language for the problem statement” (Newell & Ernst, 1965); and (8) having a “general” internal representation (Newell & Ernst, 1965; McCarthy & Hayes, 1981).
As Paolo et al. (2024) note, despite the multiple possible meanings of “generality”, most papers do not define generality even if it’s central to their argument. Without formal definition, assessing or improving generalization becomes challenging. Assuming that “generalization” is desirable while acknowledging its poor definition is misguided. We should first define specific, measurable properties before arguing that they are desirable.
Without proper definition, the value of generality remains unclear. Different types of generality support different future visions, raising unexplored questions about their relative importance. Vague definitions of generality lead to bad science and engineering (§2.2). For example, Altmeyer et al. (2024) note how the pursuit of generality has led to vague task specifications. In parallel, Gebru & Torres (2024) argue that some conceptions of AGI contravene good engineering practices: it is hard to test the functionality of systems under “standard operating conditions” if the system is advertised as a “universal algorithm for learning and acting in any environment”.
Aiming to achieve certain forms of generality could also mean making a tradeoff with ecological validity, as argued by Saxon et al. (2024). They argue that “holistic” benchmarks tend to be in practice a collection of disparate specific benchmark tasks, meaning that they have task-level construct validity. However, these tasks do not always match with user-relevant capabilities.
@@ 2.6 Normalized Exclusion
Excluding communities and experts from shaping the goals of AI research limits development of innovative and impactful ideas.
@@
Problem 1: Excluding communities. Many communities are left out of meaningful participation in shaping the goals of AI research (Delgado et al., 2023). Excluding communities causes serious harm, especially to minoritized communities (Pierre et al., 2021); it also undermines the utility of end products, reduces model performance, oversimplifies technical challenges (Kierans et al., 2025), and impedes innovation (e.g., Burt, 2004). For instance, the infamous case of facial recognition engines—such as those used by Google, Apple, and Meta—mistaking Black people for gorillas (BBC News, 2015) is still occurring more than 8 years after the problem was first identified (Grant & Hill, 2023; Appelman, 2023), with downstream impacts on surveillance and law enforcement (Jones, 2020; Pour, 2023). Similarly, selective forms of inclusion in data annotation raise ethical and practical concerns about downstream effects (Wang et al., 2022; Bertelsen et al., 2024). Other work on exclusion or inclusion of communities impacting performance is plentiful (see Buolamwini & Gebru, 2018; Young et al., 2019; Raji et al., 2020; Andrews et al., 2024; Salavati et al., 2024; Bergman et al., 2024; Weidinger et al., 2024). Excluding communities from meaningful feedback also undermines societal goals, like fostering collective legitimacy via accountability to impacted communities (Birhane et al., 2022a; Young et al., 2024; Costanza-Chock, 2020; Mikesell et al., 2013; Schulz et al., 2002).
@@
The recent prominence of AGI discourse intensifies the existing problem of community exclusion in AI research (Kelly, 2024; Frank et al., 2017). Many proponents of AGI envision a future where AI systems perform an extraordinary range of tasks for countless communities. However, research and design processes fall short of the inclusiveness demanded by this ambitious vision. For example, December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (Zeff, 2025; OpenAI, 2025), a stark departure from OpenAI’s public definition of AGI (OpenAI, 2018). As several authors argue, economic value is not the only type of desirable value (Morris et al., 2023; Pierson et al., 2025; Dulka, 2022; Harrigian et al., 2023; Agrawal et al., 2022). The economic definition helps guide the engineering decisions of OpenAI. But it is questionable whether an emphasis on profits will lead to the most beneficial or useful end products or to meaningful consensus about goals, especially for minoritized groups.
@@
Problem 2: Excluding disciplines. From application domains (e.g. medicine (Obermeyer et al., 2019), finance (Cao, 2022), cybersecurity (Salem et al., 2024), learning (Leong & Linzen, 2024)) to the practices involved in building AI—data annotation, qualitative and quantitative methods, domain expertise, computer science, and many more (Wang et al., 2022; Bertelsen et al., 2024; Widder, 2024)—AI research crosses disciplinary boundaries. The cross-disciplinary challenges of AI research mirror those of other disciplines (Stirling, 2014; Amoo et al., 2020; Stokols et al., 2003; Vestal & Mesmer-Magnus, 2020; The Royal Society & Leontidis, 2024).
@@
One major challenge is disciplinary silos, where knowledge is inadequately shared across disciplines (Ballantyne, 2019; DiPaolo, 2022; The Royal Society & Leontidis, 2024; Stirling, 2014; Amoo et al., 2020; Stokols et al., 2003). For instance, lack of knowledge sharing could be partly responsible for low attention to the distinction between explanatory and exploratory research in ML, as discussed in §2.2 (Herrmann et al., 2024).
@@
Another challenge is epistemic hierarchies—where the expertise of some disciplines is explicitly or implicitly devalued (Knorr Cetina, 1999, 2007; Fourcade et al., 2015; Simonton, 2004; Graziul et al., 2023). This can manifest as expert groups being limited to narrow input rather than contributing to broader research design decisions (Bertelsen et al., 2024).
@@
AI researchers’ focus on applying their work to other domains creates another major challenge. Insufficient domain knowledge might affect the functionality of AI tools—whether they operate as advertised (Raji et al., 2022). For instance, AI tools are deployed to make predictions about future individual-level outcomes, from pre-trial risk prediction and predictive policing to automated employment decisions (Wang et al., 2024). Yet inadequate evidence of effectiveness often fails to prevent predictive tools from being built, marketed, and deployed (Connealy et al., 2024; Doucette et al., 2021; Cameron, 2023).
@@
The problem of disciplinary silos becomes particularly acute in AGI-oriented research due to two factors: its claims to be creating cognates or replacements of human intelligence, and its claims to expertise in many disciplines. In the former case, claims are often made while ignoring debates in cognitive science and psychology about the nature of intelligence (Mitchell, 2024; Guest & Martin, 2024; Van Rooij et al., 2024; Summerfield, 2023). In the latter case, achieving AGI is often framed in terms of being able to “replace” domain experts in various domains—which are then often taken up uncritically by the media without input from domain experts themselves (e.g., Henshall, 2024).
@@
Problem 3: Resource disparities. In recent years, we have witnessed an unprecedented growth in computational resources required for training machine learning models, with requirements doubling approximately every few months (Sevilla et al., 2022). This trend compounds existing resource disparities, as state-of-the-art AI research often relies on access to computational resources accessible to very few researchers (Yu et al., 2023; LaForge, 2024). The financial cost of these resources excludes a wide range of actors from contributing to AI research. Even top research universities have a fraction of the computational resources that many corporations use to advance AI research. Efforts are underway to address this resource disparity by supporting access to large-scale computational resources maintained by government entities (e.g., NAIRR in the United States). Yet, resource parity is an aspirational goal in response to widespread recognition that AI researchers in industry enjoy a de facto advantage in setting the goals of AI research due to their access to industrial scale computational resources. This structural advantage is reinforced by the use of pre-print archives to publicize AI research without peer review (Devlin et al., 2019; Vaswani et al., 2023; Rombach et al., 2022; OpenAI et al., 2024; Bubeck et al., 2023), a strategy which legitimizes this work as scientific in nature without applying traditional standards for scientific integrity (Tenopir et al., 2016; Lin et al., 2020; Soderberg et al., 2020; Kwon & Porter, 2025; Rastogi et al., 2022).
@@
While resource disparities exist for all forms of AI research, they are particularly stark when AGI is taken as a north-star goal for the discipline, due to the orientation of current AGI efforts towards sheer computational scale and the concentration of such efforts in large tech companies. [^5] Large-scale efforts also have detrimental impacts on climate, reinforcing resource disparities further (Kaack et al., 2022; Bucknall & Dori-Hacohen, 2022; Luccioni et al., 2024). Concentration of power makes it even more important that those efforts include, rather than exclude, relevant communities and experts. AGI discourse accelerates the existing trend in AI of discounting domain expertise and lived experiences in favor of models that are allegedly experts in everything.
@@ 3 Recommendations
We have argued that AGI discourse hinders setting well-motivated science and engineering goals in AI development, while being destructive to the development of AI that has social merit. We now provide three recommendations for avoiding these traps.
@@
More specific definitions of tangible scientific, engineering, and societal goals promote a shared understanding of these goals, and thus, the capacity to evaluate whether these goals are well-motivated. Without such specificity, researchers, practitioners, and others can develop divergent understandings of a goal and how it should be achieved. This divergence enables conceptual arbitrage on the part of AI researchers and practitioners who seek to advance their own goals, since these actors ultimately determine the details of model development and implementation. People external to tech development may then be left assuming that a system achieves one goal when, in fact, it does not.
This recommendation addresses the Illusion of Consensus trap (§2.1), promoting conceptual clarity as an essential part of goal-setting. Clarity also helps to avoid the Goal Lottery trap (§2.4) by making goal selection explicit. It similarly addresses the Presuming Value-Neutrality Trap (§2.3) by explicitly surfacing values tied to specific goals, and directly reduces Generality Debt (§2.5) . Finally, goal specificity addresses the underspecification issues highlighted in the Supercharging Bad Science Trap (§2.2).
@@
In practice, pluralism can manifest in how resources are distributed among different research approaches. For example, rather than investing most computational resources in pursuit of AGI, they could be more evenly distributed among diverse goals and approaches within AI. Researchers who study the dynamics of knowledge production and problem-solving in groups have found pluralism to be beneficial (Hong & Page, 2004; Muldoon, 2013), including unique benefits ascribable to egalitarian group dynamics (Xu et al., 2022). Similarly, researchers have found that methodological diversity is beneficial to advancing knowledge (Midgley, 2000; Veit, 2020; Zhu, 2022).
@@
Inclusion supports innovation (Zhang et al., 2021; Xu et al., 2022; Burt, 2004; Hewlett et al., 2013). Identifying worthwhile goals, related use cases, and potential unintended consequences depends on engaging diverse viewpoints. Within AI research, these viewpoints must include those of end users, experts from other fields, those affected by research outcomes, and data annotators. Excluding any of these groups impoverishes the potential of AI to achieve worthwhile goals since it would discount the perspectives that define these goals as worthwhile. Including them enriches AI research.
Cross-pollination of ideas between disciplines leads to more impactful research (Shi & Evans, 2023; Dori-Hacohen et al., 2021). Such impact requires that we abandon silos of (tacit) knowledge (e.g., epistemic cultures: Knorr Cetina, 1999, 2007) and prioritize epistemic hierarchies that value non-computational research (Fourcade et al., 2015; Simonton, 2004). While technical complexity can make participation by non-experts challenging (Pierre et al., 2021), working through these challenges can surface issues experts have not anticipated (Cooper et al., 2022) and enable practical scientific contributions by integrating the insights and experiences of non-experts, or experts in other fields, into system design decisions (Delgado et al., 2023; Salavati et al., 2024).
@@
This recommendation addresses the Normalized Exclusion trap (§2.6) by treating inclusion as essential to innovation. Moreover, it addresses the Illusion of Consensus and the Presuming Value-Neutrality traps (§2.12.3) by acknowledging socially significant disagreements about the value-laden goals of AI research and development and using those disagreements productively.
4 Conclusion

### v4

@@ 2.5 Generality Debt
Relying on the generality or flexibility of tools to postpone crucial engineering, scientific, or societal decisions
@@
This trap includes the appeal to many different notions of generality at play in machine learning: (1) variety of tasks (Hernández-Orallo & Seán Ó hÉigeartaigh, 2018); (2) capability to be trained for “any task” vs. ability to perform many predefined tasks (Hernández-Orallo & Seán Ó hÉigeartaigh, 2018); (3) whether the task or data distribution the model is being evaluated on is “seen” or “unseen” (i.e., available, or not, to the model during its training phase) (Altmeyer et al., 2024); (4) variety of data in model input/output, such as structured vs unstructured, modality, etc.; (5) whether the performance of the model reflects “performance considered ‘surprising’ to humans” (Altmeyer et al., 2024); (6) variety of goals; (7) ability to “accept a general language for the problem statement” (Newell & Ernst, 1965); and (8) having a “general” internal representation (Newell & Ernst, 1965; McCarthy & Hayes, 1981).
As Paolo et al. (2024) note, despite the multiple possible meanings of “generality”, most papers do not define generality even if it is central to their argument. Without formal definition, assessing or improving generalization becomes challenging. Assuming that “generalization” is desirable while acknowledging its poor definition is misguided. We should first define specific, measurable properties before arguing that they are desirable.
Without proper definition, the value of generality remains unclear. Different types of generality support different future visions, raising unexplored questions about their relative importance. Further, vague definitions of generality lead to bad science and engineering (§2.2). For example, Altmeyer et al. (2024) note how the pursuit of generality has led to vague task specifications. In parallel, Gebru & Torres (2024) argue that some conceptions of AGI contravene good engineering practices: it is hard to test the functionality of systems under “standard operating conditions” if the system is advertised as a “universal algorithm for learning and acting in any environment.”
Aiming to achieve certain forms of generality could also mean making a tradeoff with ecological validity, as argued by Saxon et al. (2024). They argue that, in practice, “holistic” benchmarks tend to be a collection of disparate specific benchmark tasks, meaning that they have task-level construct validity. However, these tasks do not always match with user-relevant capabilities. Methodological challenges to achieving such capabilities may or may not be overcome, in time, given innovative solutions, but the pursuit of AGI assumes such challenges are surmountable. Methodological issues often inspire novel solutions, but we cannot assume a solution will be found, nor can those pursuing AGI. Further, strategies for pursuing AGI may introduce or reveal new challenges to achieving user-relevant capabilities.
@@ 2.6 Normalized Exclusion
Excluding communities and experts from shaping goals
@@
Problem 1: Excluding communities. Many communities are left out of meaningful participation in shaping the goals of AI research (Delgado et al., 2023). Excluding communities causes serious harm, especially to minoritized communities (Pierre et al., 2021); it also undermines the utility of end products, reduces model performance, oversimplifies technical challenges (Kierans et al., 2025), and can impede innovation (e.g., Burt, 2004). For instance, the infamous case of facial recognition engines—such as those used by Google, Apple, and Meta—mistaking Black people for gorillas (BBC News, 2015) is still occurring more than 8 years after the problem was first identified (Appelman, 2023; Grant & Hill, 2023), with downstream impacts on surveillance and law enforcement (Jones, 2020; Pour, 2023). Similarly, selective forms of inclusion in data annotation raise ethical and practical concerns about downstream effects (Wang et al., 2022; Bertelsen et al., 2024). Other examples of exclusion or inclusion of communities impacting performance are plentiful (see Buolamwini & Gebru, 2018; Young et al., 2019; Raji et al., 2020; Andrews et al., 2024; Bergman et al., 2024; Salavati et al., 2024; Weidinger et al., 2024). Excluding communities from meaningful feedback also undermines societal goals, such as fostering collective legitimacy through accountability to impacted communities (Schulz et al., 2002; Mikesell et al., 2013; Costanza-Chock, 2020; Birhane et al., 2022a; Young et al., 2024).
@@
The recent prominence of AGI discourse intensifies the existing problem of community exclusion in AI research (Frank et al., 2017; Kelly, 2024). Many proponents of AGI envision a future where AI systems perform an extraordinary range of tasks for countless communities. However, research and design processes fall short of the inclusiveness demanded by this ambitious vision. For example, December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (OpenAI, 2025d; Zeff, 2025), a stark departure from OpenAI’s public definition of AGI (OpenAI, 2018). As several authors argue, economic value is not the only type of desirable value (Agrawal et al., 2022; Dulka, 2022; Harrigian et al., 2023; Morris et al., 2024; Pierson et al., 2025). The economic definition helps guide the engineering decisions of OpenAI. However, it is questionable whether an emphasis on profits will lead to the most beneficial or useful end products or to meaningful consensus about goals, especially for minoritized groups.
@@
Problem 2: Excluding disciplines. From application domains (e.g., medicine (Obermeyer et al., 2019), finance (Cao, 2022), cybersecurity (Salem et al., 2024), learning (Leong & Linzen, 2024)) to the practices involved in building AI—data annotation, qualitative and quantitative methods, domain expertise, computer science, and many more (Wang et al., 2022; Bertelsen et al., 2024; Widder, 2024)—AI research crosses disciplinary boundaries. The cross-disciplinary challenges of AI research mirror those of other disciplines (Stokols et al., 2003; Stirling, 2014; Amoo et al., 2020; Vestal & Mesmer-Magnus, 2020; The Royal Society, 2024).
@@
One major challenge is disciplinary silos, where knowledge is inadequately shared across disciplines (Stokols et al., 2003; Stirling, 2014; Ballantyne, 2019; Amoo et al., 2020; DiPaolo, 2022; The Royal Society, 2024). For instance, lack of knowledge sharing could be partly responsible for low attention to the distinction between explanatory and exploratory research in ML, discussed in §2.2 (Herrmann et al., 2024).
@@
Another challenge is epistemic hierarchies—where the expertise of some disciplines is explicitly or implicitly devalued (Knorr Cetina, 1999, 2007; Simonton, 2004; Fourcade et al., 2015; Graziul et al., 2023). This can manifest as expert groups being limited to narrow input rather than contributing to broader research design decisions (Bertelsen et al., 2024).
@@
AI researchers’ focus on applying their work to other domains creates another major challenge. Insufficient domain knowledge might affect the functionality of AI tools—whether they operate as advertised (Raji et al., 2022). For instance, AI tools are deployed to make predictions about future individual-level outcomes, from pre-trial risk prediction and predictive policing to automated employment decisions (Wang et al., 2024). Yet inadequate evidence of effectiveness often fails to prevent predictive tools from being built, marketed, and deployed (Doucette et al., 2021; Cameron, 2023; Connealy et al., 2024).
@@
The problem of disciplinary silos becomes particularly acute in AGI-oriented research due to two factors: its claims to be creating cognates or replacements of human intelligence, and its claims to expertise in many disciplines. In the former case, claims are often made while ignoring debates in cognitive science and psychology about the nature of intelligence (Summerfield, 2023; Guest & Martin, 2024; Mitchell, 2024; Van Rooij et al., 2024). In the latter case, achieving AGI is often framed in terms of being able to “replace” domain experts in various domains—which are then often taken up uncritically by the media without input from domain experts themselves (e.g., Henshall, 2024).
@@
Problem 3: Resource disparities. In recent years, we have witnessed an unprecedented growth in computational resources required for model training, with requirements doubling approximately every few months (Sevilla et al., 2022). This trend compounds existing resource disparities, as state-of-the-art AI research often relies on access to computational resources accessible to very few researchers (Yu et al., 2023; LaForge, 2024). The financial cost of these resources excludes a wide range of actors from contributing to AI research, as even top research universities have a fraction of the computational resources that many corporations use to advance AI research. Efforts are underway to address this resource disparity by supporting access to large-scale computational resources maintained by government entities (e.g., NAIRR in the United States). Yet, resource parity is an aspirational goal in response to widespread recognition that AI researchers in industry enjoy a de facto advantage in setting the goals of AI research due to their access to industrial scale computational resources. This structural advantage is reinforced by the use of pre-print archives to publicize AI research without peer review (Devlin et al., 2019; Rombach et al., 2022; Bubeck et al., 2023; OpenAI et al., 2024), a strategy which legitimizes this work as scientific in nature without applying traditional standards for scientific integrity (Tenopir et al., 2016; Lin et al., 2020; Soderberg et al., 2020; Rastogi et al., 2022; Kwon & Porter, 2025).
@@
While resource disparities exist for all forms of AI research, they are particularly stark when AGI is taken as a north-star goal for the discipline, due to the orientation of current AGI efforts towards sheer computational scale and the concentration of such efforts in large tech companies. [^7] Large-scale efforts also have detrimental impacts on climate, reinforcing resource disparities (Bucknall & Dori-Hacohen, 2022; Kaack et al., 2022; Luccioni et al., 2024). Such concentration of power makes it even more important that those efforts include, rather than exclude, relevant communities and experts. That is, AGI discourse accelerates the existing trend in AI of discounting domain expertise and lived experiences in favor of models that are allegedly experts in everything.
@@ 3 Recommendations
We have argued that AGI discourse hinders setting well-motivated scientific and engineering goals in AI development, while being destructive to the development of AI that has social merit. We now provide three recommendations for avoiding these traps.
@@
More specific definitions of tangible scientific, engineering, and societal goals promote a shared understanding of these goals, and thus the capacity to evaluate whether these goals are well-motivated. Without such specificity, researchers, practitioners, and others can develop divergent understandings of a goal and how it should be achieved. This divergence enables conceptual arbitrage on the part of AI researchers and practitioners who seek to advance their own goals, since these actors ultimately determine the details of model development and implementation. People external to AI development may then be left assuming that a system achieves a specific goal when, in fact, it does not.
Specificity can maintain sufficient flexibility for exploratory research by engaging in best practices around developing a research question. Consider the research goal “How can a mixture of experts (MoE) strategy improve performance of Whisper large-v3 in challenging speech domains?” which could cover various domains of speech or MoE strategies. A reformulated, more specific goal would be: “How can a MoE strategy, where experts are a series of models fine-tuned on short (<<<2s), medium (2–20s), and long utterances (>>>20s), help improve performance of Whisper large-v3 in a speech domain dominated by short utterances but also containing relatively long utterances?” Without additional specification, answering the first question provides little guarantee that a solution would address the specific features of the second question, at least not without substantial effort to understand how such a general solution may be applied to this specific case/domain. This example is illustrative, though based on real features of a speech domain where accuracy is essential (i.e., police radio communications, see Srivastava et al. 2024; Venkit et al. 2024).
Goal Specificity addresses the Illusion of Consensus trap (§2.1), promoting conceptual clarity as an essential part of goal-setting. Clarity also helps to avoid the Goal Lottery trap (§2.4) by making goal selection explicit. It similarly addresses the Presuming Value-Neutrality Trap (§2.3) by explicitly surfacing values tied to specific goals, and directly reduces Generality Debt (§2.5) . Finally, goal specificity addresses the underspecification issues highlighted in the Supercharging Bad Science Trap (§2.2).
@@
Researchers who study the dynamics of knowledge production and problem-solving in groups have found pluralism to be beneficial (Hong & Page, 2004; Muldoon, 2013), including unique benefits ascribable to egalitarian group dynamics (Xu et al., 2022).
Pluralism in AI research can take several forms, each with distinct implications for how we approach complex problems. For example:
Methodological pluralism implies that different ways of approaching a problem help achieve better solutions, often an effective strategy for complex problem-solving in general (Midgley, 2000; Veit, 2020; Zhu, 2022) .
Value pluralism, as applied to alignment research, implies a direct connection between technical advancement and accommodation of plural values (Sorensen et al., 2024a, b).
Algorithmic pluralism addresses the “patterned inequality” associated with algorithmic monoculture and implies a “plurality of paths to different outcomes” must be supported to avoid reproducing existing social inequalities (e.g., embedded in data) (Jain et al., 2024) .
Each form of pluralism translates into concrete research practices, such as choice of method(s), setting of pluralistic goals, and testing algorithms to ensure known harms (i.e., algorithmic discrimination) are addressed.
To name just one example, algorithmic decision-making in hiring processes is unlikely to benefit from AGI as much as from targeted solutions to that particular setting, which account for the domain-specific nature of most positions, the different ways hiring managers evaluate candidates, and existing evidence of algorithmic discrimination in hiring. In practice, pluralism can manifest in how resources are distributed among different research approaches. For example, rather than investing most computational resources in pursuit of AGI, they could be more evenly distributed among diverse goals and approaches within AI.
@@
Inclusion supports innovation (Burt, 2004; Hewlett et al., 2013; Zhang et al., 2021; Xu et al., 2022). Identifying worthwhile goals, related use cases, and potential unintended consequences depends on engaging diverse viewpoints. Within AI research, these viewpoints must include those of end users, experts from other fields, those affected by research outcomes, and data annotators. Excluding any of these groups impoverishes the potential of AI to achieve worthwhile goals since it would discount the perspectives that define these goals as worthwhile. Including them enriches AI research.
Cross-pollination of ideas between disciplines leads to more impactful research (Dori-Hacohen et al., 2021; Shi & Evans, 2023). Such impact requires that we abandon silos of (tacit) knowledge (e.g., epistemic cultures: Knorr Cetina, 1999, 2007) and prioritize epistemic hierarchies that value non-computational research (Simonton, 2004; Fourcade et al., 2015). While technical complexity can make participation by non-experts challenging (Pierre et al., 2021), working through these challenges can surface issues experts have not anticipated (Cooper et al., 2022) and enable practical scientific contributions by integrating the insights and experiences of non-experts, or experts in other fields, into system design decisions (Delgado et al., 2023; Salavati et al., 2024).
@@
This recommendation addresses the Normalized Exclusion trap (§2.6) by treating inclusion as essential to innovation. Moreover, it addresses the Illusion of Consensus and the Presuming Value-Neutrality traps (§2.1, §2.3) by acknowledging socially significant disagreements about the value-laden goals of AI research and development and using those disagreements productively.
4 Alternative Views